# Plasmonic metamaterial time crystal

### v1

@@ Discussion
We obtain a maximum modulation parameter $\eta_{0}/\omega_{\mathrm{c}}^{0}=\max\limits_{\tau}|\eta(\tau)|/\omega_{\mathrm{c}}^{0}=0.18\pm 0.02$ for which the model reproduces the experimental data with very good accuracy. This corresponds to an impressive $\delta m^{*}(\tau)/m^{*}_{0}\approx 80\%$ dynamical effective mass modulations. As a function of driving strength, we further establish that these parametric effects are saturated for the reported peak driving field of 40kV/cm. This is indicated by the leveling-off of the response and the emergence of incoherent nonlinearities measured at a larger field of $140\textrm{kV/cm}$ (see supplementary material). Hence, this value appears to be a physical bound for coherent effective mass modulations in this system. Given this value, a strategy to enhance further parametric effects consists in increasing the ratio $L_{K}^{0}/L_{G}$. It can be achieved by decreasing the size of the plasmonic cavities so as to operate in the ’electrostatic’ regime where retardation effects vanish and kinetic ones dominate Staffaroni_2012 ; Greffet_2012 : $L_{K}^{0}\gg L_{G}$ and $\beta=1$. In the following, using Floquet analysis with periodic drive $\eta(\tau)=-\eta_{0}\cos^{2}(\omega_{\mathrm{d}}\tau)$ and accounting for the observed losses, we highlight three key physical phenomena lying within the experimentally accessible range of parameters of this system for : 1) the present cavity design ($\beta\approx 0.65,L_{K}^{0}/L_{G}\approx 2.4$, $\eta_{0}/\omega_{\mathrm{c}}^{0}\approx 0.18$) and 2) an electrostatic plasmonic cavity ($\beta=1,L_{K}^{0}\gg L_{G}$, $\eta_{0}/\omega_{\mathrm{c}}^{0}=0.4$).
Parametric amplification of surface plasmons
As well known from other parametrically driven systems, the modulation induced by the drive generates gain on the cavity mode. The corresponding growth rate of the mode $\mathrm{Im}(\mu)/\omega_{\mathrm{c}}^{0}$, with $\mu$ being an eigenvalue of the Floquet matrix describing long-term behavior of the system (see methods), is reported in the phase diagram of Fig. 3a as a function of drive frequency and modulation strength, together with the operating points of the two aforementioned cavity designs. When the growth rate overcomes the total loss rate $\Gamma/\omega_{\mathrm{c}}^{0}$ (including radiative and non-radiative contributions), the system develops a parametric instability: it crosses a threshold illustrated by the solid contour line and defined by $\mathrm{Im}(\mu)/\omega_{\mathrm{c}}^{0}=\Gamma/\omega_{\mathrm{c}}^{0}$ in Fig. 3a for which quantities diverge exponentially and parametric amplification sets in. Even before reaching this threshold and within the broader region defined by the dashed contour line, the system remains stable as a function of time but demonstrates average reflectivities above 1 (see methods and supplementary material). This highlights the strong potential of this platform for achieving light amplification and lasing of surface plasmons.
Figure 3: Plasmonic time crystal phenomena computed for the experimentally accessible regime of parameters of this platform
a) Stability diagram: growth rate of the cavity mode as a function of drive frequency $\omega_{\mathrm{d}}$ and modulation strength $\eta_{0}$ relative to the equilibrium cavity frequency $\omega_{\mathrm{c}}^{0}$. The solid contour line defined by $\mathrm{Im}(\mu)/\omega_{\mathrm{c}}^{0}=\Gamma/\omega_{\mathrm{c}}^{0}=0.12$ marks the threshold for parametric amplification given the observed losses. Dashed contour line indicate regions exhibiting average reflectivity above one. Operating points of the present cavity design (star) and an electrostatic cavity (dot) are indicated (see text) b) Wigner function of the cavity state accounting for the losses. Top panel: no drive. Middle panel: this experiment’s parameters ($\eta_{0}/\omega_{\mathrm{c}}^{0}=0.18$, $\omega_{\rm d}/\omega_{\mathrm{c}}^{0}=0.92$) Bottom panel: an electrostatic cavity ($\eta_{0}/\omega_{\mathrm{c}}^{0}=0.4$, $\omega_{\rm d}/\omega_{\mathrm{c}}^{0}=0.8$) c Plasmonic time crystal $k$-gap openings. Real and imaginary parts of the eigenmodes’ frequency $\omega$ as a function of momentum (or Bloch-momentum) $k$, normalized to the plasma frequency $\omega_{\mathrm{p}}$ and including the losses, for the equilibrium (dashed lines) and the temporally-modulated systems (solid lines). Top panel: a dispersive insulator/plasmonic interface. Bottom panel: our dispersion-free plasmonic metamaterial design for cavities tuned in the electrostatic regime (shown in the first Brillouin zone of the metamaterial of period $d$. The drive frequency corresponds to $\omega_{\mathrm{d}}/\omega_{\mathrm{p}}=0.68$ in both cases.
Generation of squeezed and entangled plasmons
[…]
$k$-agnostic gap openings in plasmonic metamaterial time crystals
Moving from a single cavity picture to the actual periodically structured metamaterial, a central aspect of time crystals lies in the opening of $k$-gaps in their dispersion relations Asgari_2024 , where $k$ denotes the momentum (translational invariant systems) or quasi-momentum (periodically structured metamaterials) of photonic eigenmodes. Since material dispersion limits the width of the $k$-gaps, several strategies were recently proposed to widen their span Wang_2024 ; Feinberg_2025 . The ability of our metamaterial approach to engineer a flat dispersion allows to address such challenges Aupiais_2023 . In Fig 3c. we compare a dispersive insulator/plasmonic interface together with our metamaterial design and compute $k$-gap openings in the dispersion relation (see methods). For the dispersive plasmonic interface (top panel), although the $k$-gaps are already significant, they fail to cover the whole dispersion. In contrast, our dispersion-free metamaterial (bottom panel) achieves $k$-gap openings across the full Brillouin zone, effectively realizing a $k$-agnostic time crystal.
@@ Conclusion
In conclusion, we report the first optical implementation of a photonic time crystal. This is realized in a plasmonic system and enabled by the sizeable and coherent dynamical variations of the carriers’ effective mass. Our work establishes that this mechanism, previously known in bulk systems, can be extended to surface plasmons and leveraged as a foundational element for constructing plasmonic metamaterials harbouring Floquet physics and time crystalline phenomena. Along those lines, the inherent tunability of the plasmonic medium to external parameters Aupiais_2023 ; Aupiais_2024 offers a unique route towards extension of PTC phenomena to other photonic degrees of freedom, including spin and orbital ones.
@@ Methods
A $\langle 100\rangle$-oriented, nominally undoped InSb crystal with dimensions of $5\penalty 10000\ \mathrm{mm}\times 5\penalty 10000\ \mathrm{mm}\times 0.5\penalty 10000\ \mathrm{mm}$ is used as the starting material. A $\sim 3\penalty 10000\ \mu\mathrm{m}$-thick $\mathrm{Si}_{3}\mathrm{N}_{4}$ insulating layer is first deposited via plasma-enhanced chemical vapor deposition. The top metallic pattern is then defined through photolithography, followed by Ti/Au (20 nm/200 nm) evaporation and lift-off. The resulting metamaterial structure consists of periodic metallic stripes with a width of $s=41\penalty 10000\ \mu\mathrm{m}$ and spacing $a=16\penalty 10000\ \mu\mathrm{m}$, yielding a total period of $d=s+a=57\penalty 10000\ \mu\mathrm{m}$ and follows the design reported in a previous work Aupiais_2023 .
@@
The inductance $L$ has two contributions Staffaroni_2012 :
@@
Computation of the parameters entering equations (8) and (9) is outlined here and explained in more detail in the supplementary material. These calculations are based on the determination of the cavity resonance frequency $\omega_{\rm c}$, obtained by solving the dispersion relation of a truncated metal/insulator/plasmonic waveguide Aupiais_2023 , when parameters are varied. Notably, the resonance frequency depends, among other geometrical and optical parameters of the constitutive elements of the metamaterial, on the plasma frequency of the plasmonic medium supporting the resonance: $\omega_{\rm p}=\sqrt{\frac{ne^{2}}{\epsilon_{0}\epsilon_{\infty}m^{*}}}$, where $n$ and $e$ are the density and charge of the carriers, $m^{*}$ their effective mass, $\epsilon_{0}$ the vacuum permittivity and $\epsilon_{\infty}$ the high-frequency permittivity of the plasmonic medium. Therefore, the ratio of kinetic to geometric inductance $L_{K}^{0}/L_{G}$ at equilibrium can be determined from the resonance frequencies $\omega_{\mathrm{c}}^{0}$ and $\omega_{0}$ obtained by varying the plasma frequency for two otherwise identical cavities: 1) with equilibrium plasma frequency, yielding $\omega_{\rm c}=\frac{1}{\sqrt{(L_{K}^{0}+L_{G})C}}$ and 2) with infinite plasma frequency, for which the kinetic inductance vanishes while the geometric inductance and capacitance remain unchanged, yielding $\omega_{0}=\frac{1}{\sqrt{L_{G}C}}$. Similarly, the dependence of kinetic inductance with the effective mass, allowing to determine $\delta L_{K}$ and $\beta$, is obtained through a variation of the effective mass that enters the plasma frequency, all other parameters being fixed. We point out that these computations are parameter-free and obtained from experimental measurements of the geometrical and optical properties of the constitutive elements of the metamaterial, as detailed in Aupiais_2023 .
@@
To compute the temporal response of the system to the experimentally measured incident driving field $E_{d}(\tau)$ shown in Fig. 2, we determine $\eta(\tau)$ from equation (9) as follows. We consider a non-parabolic band structure given by the Kane model suitable for the description of small-band gap semiconductors Kane_1957 where the electronic dispersion relation $\varepsilon(p)$ satisfies $\varepsilon(p)(1+\alpha\varepsilon(p))=\frac{p^{2}}{2m^{*}_{0}}$ and $\alpha$ is the parameter that accounts for non-parabolicity. The effective mass $m^{*}(p)$ is defined by $m^{*}(p)=p\left(\frac{d\varepsilon}{dp}\right)^{-1}$, which allows to express it as: $m^{*}(p)=m^{*}_{0}(1+2\alpha\varepsilon(p))$ Zawadzki_1974 . Keeping only the leading order term $\propto p^{2}$ in the electronic dispersion relation, we parametrize the effective mass as $m^{*}(p)=m^{*}_{0}(1+2\alpha\frac{p^{2}}{2m^{*}_{0}})$, so that $\frac{\delta m^{*}(p)}{m^{*}_{0}}=\frac{\alpha p^{2}}{m^{*}_{0}}$. We solve the dynamics of the electrons’ momentum $p(\tau)$ in the resonator according to:
@@
Non-parabolicity parameter $\alpha$ $5.6\penalty 10000\ \mathrm{eV}^{-1}$ Zawadzki_1974
@@
Floquet analysis (Figure 3)
For the Floquet analysis presented in Fig. 3, we use a periodic drive: $\eta(\tau)=-\eta_{0}\cos^{2}(\omega_{\mathrm{d}}\tau)$.
Phase diagram (Figure 3a)
Under a periodic drive $\eta(t)$, the evolution operator of the cavity mode (a^,a^)($\hat$$^{\dagger},\hat{a}$) is decomposed using Floquet’s theorem as:

### v2

@@ Discussion
We obtain a maximum modulation parameter $\eta_{0}/\omega_{\mathrm{c}}^{0}=\max\limits_{\tau}|\eta(\tau)|/\omega_{\mathrm{c}}^{0}=0.18\pm 0.02$ for which the model reproduces the experimental data with very good accuracy. This corresponds to an impressive $\delta m^{*}(\tau)/m^{*}_{0}\approx 80\%$ dynamical effective mass modulations. This value is not limited by the maximum driving field that can be applied, which could be further increased, but rather by the onset of incoherent nonlinearities, namely impact ionization (see Supplementary Material), which starts to emerge at fields exceeding $\sim 40kV/cm$ due to the small band-gap of the InSb plasmonic material Houver et al. (2019); Biasco et al. (2022). Hence, it appears to be a physical bound for coherent effective mass modulations in this system.
Spectroscopy of a PTC
We next investigate the periodically driven quasi–steady-state regime observed at $\tau_{0}\simeq 24\,\mathrm{ps}$ by analyzing the dynamics over one period of the drive $\tau\in[\tau_{0},\tau_{0}+2\pi/\omega_{d}]$ within the framework of Floquet theory. For a system that is periodic in time with frequency $2\omega_{d}$, the complex reflectivity $\tilde{r}(\tau,\omega)$ can be expanded in Fourier harmonics at frequencies $\pm 2n\omega_{d}$ (see Supplementary material). In the present case, this expansion is dominated by the zeroth-order (time-averaged) component and the first-order harmonic contributions, such that:
$\tilde{r}(\tau,\omega)=\tilde{r}_{b}(\omega)+\tilde{r}_{0}(\omega)+\sum_{n=\pm 1}\tilde{r}_{n}(\omega)\,e^{2in\omega_{d}\tau}.$ (4)
The time-averaged response consists of a non-resonant background $\tilde{r}_{b}(\omega)$ and a resonant cavity contribution $\tilde{r}_{0}(\omega)$, while $\tilde{r}_{\pm 1}(\omega)$ are the first-order Floquet sidebands. The latter correspond to photons whose frequency has been either up-converted, $\tilde{r}_{1}(\omega)$, or down-converted, $\tilde{r}_{-1}(\omega)$, by $2\omega_{d}$ with respect to their incident frequency $\omega$. Spectra for $\tilde{r}_{0}(\omega)$ and $\tilde{r}_{\pm 1}(\omega)$ are presented in Fig. 3a. Although the reflectivity satisfies $\tilde{r}(\tau,\omega)=\tilde{r}(\tau,-\omega)^{*}$ at any time $\tau$, spectra of the Floquet sidebands $\tilde{r}_{\pm 1}(\omega)$ display two striking features: as opposed to $\tilde{r}_{0}(\omega)$, they are predominantly one-sided in frequency and their lineshapes deviate qualitatively from that expected of a simple resonance. Instead, they manifest characteristics of a second order resonance (or pole) emanating from the presence of two closely spaced resonances within the spectral linewidth (see below). As shown in Fig. 3b, these aspects can be interpreted as follows. In the absence of modulation, the cavity resonates for both positive ($\omega\approx\omega_{c}^{0}$) and negative ($\omega\approx-\omega_{c}^{0}$) frequencies within the spectral width $\Gamma$, which is described mathematically as two poles in the reflectivity located in the complex-frequency plane at $\tilde{\omega}_{c}^{0}=\omega_{c}^{0}-i\Gamma$ and $(\tilde{\omega}_{c}^{0})^{*}=-\omega_{c}^{0}-i\Gamma$ (Fig. 3b-panel (1)). The periodic drive generates Floquet replicas of these poles shifted by integer multiples of $\pm 2\omega_{d}$. Notably, when the drive frequency is tuned close to the cavity resonance ($\omega_{d}\approx\omega_{c}^{0}$), the first-order replica of the negative (resp. positive) frequency component becomes nearly degenerate with its original positive (resp. negative) frequency counterpart i.e., $\omega_{c}^{0}+2\omega_{d}\approx\omega_{c}^{0}$ (resp. $\omega_{c}^{0}-2\omega_{d}\approx-\omega_{c}^{0}$) (Fig. 3b-panel (2)). These two Floquet eigenmodes are described as two closely spaced poles, which can be written on the positive frequency side as:
$\tilde{\mu}_{\pm}=\tilde{\omega}_{d}\pm\Delta\tilde{\mu}$, (5)
where $\tilde{\omega}_{d}=\omega_{d}-i\Gamma$. One can show that the Floquet matrix displays PT symmetry (see Supplementary material) which, if unbroken, constrains $\Delta\tilde{\mu}$ to be real-valued Bender and Boettcher (1998); El-Ganainy et al. (2018), as it is the case for weak driving, see Fig. 3c. In this regime, photons undergo repeated scattering between the positive and negative frequency components of the resonance and find a resonant pathway to get either up or down converted. Importantly, non-resonant pathways in which, for instance, positive frequency photons attempt to get up-converted are strongly inhibited, explaining why the Floquet sidebands appear predominantly one-sided (e.g. $|\tilde{r}_{1}(+\omega)|\ll|\tilde{r}_{1}(-\omega)|$). Ultimately, the Floquet sidebands probe the joint spectral density of states, i.e. the spectral overlap $\propto\frac{1}{(\omega-\tilde{\mu}_{+})(\omega-\tilde{\mu}_{-})}$, of the two Floquet eigenmodes, which naturally accounts for their second-order pole character (see Supplementary material).
As the driving strength is increased, the coupling between those two eigenmodes gets stronger, their separation $\Delta\tilde{\mu}$ decreases and eventually vanishes, at which point they coalesce at $\tilde{\omega}_{d}$ (Fig. 3b-panel (3)). This coalescence defines an exceptional point characterized by a spontaneous breaking of PT symmetry Koutserimpas et al. (2018); Bossart and Fleury (2021) and marks the transition into the PTC regime. Beyond this point, the difference between the two Floquet eigenmodes $\Delta\tilde{\mu}$ become purely imaginary
$\tilde{\mu}_{\pm}=\tilde{\omega}_{d}\pm i\Delta\Gamma$ (6)
leading to reduced damping in one of the two modes, which is a hallmark of the PTC phase (Fig. 3b-panel (4)) Asgari et al. (2024); Feinberg et al. (2025).
In Fig. 3c-d, using Floquet analysis with periodic drive $\eta(\tau)=-\eta_{0}\cos^{2}(\omega_{\mathrm{d}}\tau)$, we compute the evolution of $\Delta\tilde{\mu}/\omega_{c}^{0}$ (Fig. 3c) and of the spectral lineshapes expected across this transition (Fig 3d), from which two important conclusions can be drawn: 1) the spectra evolve smoothly across the PTC transition and 2) the lineshapes’ characteristics of $\tilde{r}_{0}(\omega)$ and $\tilde{r}_{\pm 1}(\omega)$ (both real and imaginary parts) encode detailed information about the poles’ structure and the system’s parameters, notably the spectral locations of $\tilde{\mu}_{\pm}$ which can be obtained through fitting of the experimental lineshapes (see Methods and Supplementary material).
Figure 4: a Left panel: Temporal evolution of the Floquet eigenvalues $\tilde{\mu}_{\pm}$ (solid lines) shown with 95% confidence interval Right panel: Evolution of the Floquet eigenvalues $\tilde{\mu}_{\pm}$ as a function of drive strength $\eta_{0}/\omega_{c}^{0}$ computed from the model for parameters taken from the experiment. c) Stability diagram of the driven system: parametric gain $\Delta\Gamma$ as a function of drive frequency $\omega_{\mathrm{d}}$ and modulation strength $\eta_{0}$ relative to the equilibrium cavity frequency $\omega_{\mathrm{c}}^{0}$. The colored area where $\Delta\Gamma\neq 0$ signals the PTC regime. The solid contour line defined by $\Gamma/\omega_{\mathrm{c}}^{0}=0.12$ marks the threshold for plasmonic lasing given the measured equilibrium losses $\Gamma$.
Temporal transition to the PTC regime
Below, we unambiguously confirm the transition to the PTC regime by extracting the spectral locations of the Floquet eigenmodes as a function of time $\tau$. Treating the drive amplitude as a slowly varying function of time with respect to the fast modulations at $\pm 2\omega_{d}$, this is achieved by performing a short term Fourier Transform of the time-domain data (Fig. 2b) to extract the slowly varying components $\tilde{r}_{0}(\tau,\omega)$ and $\tilde{r}_{1}(\tau,\omega)$ and performing a fit of the lineshapes to extract the spectral location of the poles. The results are presented in Fig 4a. As a function of time, we observe the pole coalescence and the transition to the PTC regime at around $\tau\approx 21ps$ before the poles acquire different imaginary part. The experimental results are seen to be in excellent agreement with predictions of our model, notably in the magnitude of the effect obtained. Within the PTC regime, the emergence of gain is directly reflected in the reduction of the imaginary part of one of the poles. Experimentally, this corresponds to a linewidth narrowing of approximately $\Delta\Gamma\approx 40GHz$ from an initial equilibrium linewidth $\Gamma=\gamma_{r}+\gamma_{nr}$, with $\gamma_{nr}\approx 70GHz$ and $\gamma_{r}\approx 20GHz$ being the non radiative and radiative losses respectively. It indicates that the non-radiative plasmonic losses $\gamma_{nr}$ are reduced by more than 50%. To the best of our knowledge, this constitutes the first demonstration of plasmonic loss reduction induced by a parametric drive, opening a new route for plasmonic loss engineering based on a temporal approach Kiselev et al. (2024b); Feinberg et al. (2025).
Towards plasmonic lasing
Finally, we place these results in the broader context of this platform’s parameters and discuss the possibility of plasmonic lasing in this system, which may occur when the gain produced by the drive overcomes the total losses ($\Delta\Gamma>\Gamma$). To that end, we compute in Fig. 4b its phase diagram as a function of drive frequency and strength. The colored area marks the PTC phase with the emergence of gain ($\Delta\Gamma\neq 0$), while the white area corresponds to the absence of gain ($\Delta\Gamma=0$). The current experimental realization of Fig 4a are reported as colored dots on this diagram. We predict that plasmonic lasing with this platform is within experimental reach with an improved cavity design and fine tuning of the drive parameters. A strategy for this consists in increasing the ratio $L_{K}^{0}/L_{G}$ by decreasing the size of the plasmonic cavities so as to operate in the ’electrostatic’ regime where retardation effects vanish and kinetic ones dominate Staffaroni et al. (2012); Greffet (2012). For such an electrostatic cavity ($\beta=1,L_{K}^{0}\gg L_{G}$, $\eta_{0}/\omega_{\mathrm{c}}^{0}=0.4$, white dot in Fig 4b), this enhances parametric effects as compared to the present cavity design ($\beta\approx 0.65,L_{K}^{0}/L_{G}\approx 2.4$, $\eta_{0}/\omega_{\mathrm{c}}^{0}\approx 0.18$, star in Fig 4b), and may ultimately allow to cross the threshold for lasing, as indicated by the white solid line shown in Fig. 4.
@@ Conclusion
In conclusion, we report the first all-optical realization of a photonic time crystal, operating at THz frequencies and implemented with a plasmonic metamaterial. The underlying mechanism relies on large and ultrafast dynamical modulations of the carriers’ effective mass sustaining the plasmonic resonance. The transition to the PTC regime, evidenced by direct spectroscopic measurements and fully supported by a theoretical analysis based on Floquet theory, manifests through the coalescence of two Floquet-driven optical modes followed by the onset of plasmonic gain. This work paves the way towards the temporal engineering of dissipation and light-matter interactions in plasmonic systems and establishes a solid framework for the realization and spectroscopic identification of PTC phenomena in future implementations.
@@ Methods
A $\langle 100\rangle$-oriented, nominally undoped InSb crystal with dimensions of $5\penalty 10000\ \mathrm{mm}\times 5\penalty 10000\ \mathrm{mm}\times 0.5\penalty 10000\ \mathrm{mm}$ is used as the starting material. A $\sim 3\penalty 10000\ \mu\mathrm{m}$-thick $\mathrm{Si}_{3}\mathrm{N}_{4}$ insulating layer is first deposited via plasma-enhanced chemical vapor deposition. The top metallic pattern is then defined through photolithography, followed by Ti/Au (20 nm/200 nm) evaporation and lift-off. The resulting metamaterial structure consists of periodic metallic stripes with a width of $s=41\penalty 10000\ \mu\mathrm{m}$ and spacing $a=16\penalty 10000\ \mu\mathrm{m}$, yielding a total period of $d=s+a=57\penalty 10000\ \mu\mathrm{m}$ and follows the design reported in a previous work Aupiais et al. (2023).
@@
The inductance $L$ has two contributions Staffaroni et al. (2012):
@@
Computation of the parameters entering equations (11) and (12) is outlined here and explained in more detail in the supplementary material. These calculations are based on the determination of the cavity resonance frequency $\omega_{\rm c}$, obtained by solving the dispersion relation of a truncated metal/insulator/plasmonic waveguide Aupiais et al. (2023), when parameters are varied. Notably, the resonance frequency depends, among other geometrical and optical parameters of the constitutive elements of the metamaterial, on the plasma frequency of the plasmonic medium supporting the resonance: $\omega_{\rm p}=\sqrt{\frac{ne^{2}}{\epsilon_{0}\epsilon_{\infty}m^{*}}}$, where $n$ and $e$ are the density and charge of the carriers, $m^{*}$ their effective mass, $\epsilon_{0}$ the vacuum permittivity and $\epsilon_{\infty}$ the high-frequency permittivity of the plasmonic medium. Therefore, the ratio of kinetic to geometric inductance $L_{K}^{0}/L_{G}$ at equilibrium can be determined from the resonance frequencies $\omega_{\mathrm{c}}^{0}$ and $\omega_{0}$ obtained by varying the plasma frequency for two otherwise identical cavities: 1) with equilibrium plasma frequency, yielding $\omega_{\rm c}=\frac{1}{\sqrt{(L_{K}^{0}+L_{G})C}}$ and 2) with infinite plasma frequency, for which the kinetic inductance vanishes while the geometric inductance and capacitance remain unchanged, yielding $\omega_{0}=\frac{1}{\sqrt{L_{G}C}}$. Similarly, the dependence of kinetic inductance with the effective mass, allowing to determine $\delta L_{K}$ and $\beta$, is obtained through a variation of the effective mass that enters the plasma frequency, all other parameters being fixed. We point out that these computations are parameter-free and obtained from experimental measurements of the geometrical and optical properties of the constitutive elements of the metamaterial, as detailed in Aupiais et al. (2023).
@@
To compute the temporal response of the system to the experimentally measured incident driving field $E_{d}(\tau)$ shown in Fig. 2, we determine $\eta(\tau)$ from equation (12) as follows. We consider a non-parabolic band structure given by the Kane model suitable for the description of small-band gap semiconductors Kane (1957) where the electronic dispersion relation $\varepsilon(p)$ satisfies $\varepsilon(p)(1+\alpha\varepsilon(p))=\frac{p^{2}}{2m^{*}_{0}}$ and $\alpha$ is the parameter that accounts for non-parabolicity. The effective mass $m^{*}(p)$ is defined by $m^{*}(p)=p\left(\frac{d\varepsilon}{dp}\right)^{-1}$, which allows to express it as: $m^{*}(p)=m^{*}_{0}(1+2\alpha\varepsilon(p))$ Zawadzki (1974). Keeping only the leading order term $\propto p^{2}$ in the electronic dispersion relation, we parametrize the effective mass as $m^{*}(p)=m^{*}_{0}(1+2\alpha\frac{p^{2}}{2m^{*}_{0}})$, so that $\frac{\delta m^{*}(p)}{m^{*}_{0}}=\frac{\alpha p^{2}}{m^{*}_{0}}$. We solve the dynamics of the electrons’ momentum $p(\tau)$ in the resonator according to:
@@
Non-parabolicity parameter $\alpha$ $5.6\penalty 10000\ \mathrm{eV}^{-1}$ Zawadzki (1974)
@@
Spectroscopic analysis based on Floquet theory (Figures 3 & 4)
Floquet theory
Under a drive $\eta(\tau)=-\eta_{0}\cos^{2}(\omega_{d}\tau)$ which is periodic with frequency $2\omega_{d}$, the evolution operator of the cavity mode (a^,a^)($\hat$$^{\dagger},\hat{a}$) is decomposed using Floquet’s theorem as:
@@
where $\mathbf{G}(\tau)$ and $\mathbf{\mu}$ are both $2\times 2$ matrices acting on the spinor space of the cavity mode (a^,a^)($\hat$$^{\dagger},\hat{a}$). $\mathbf{G}(\tau)$ is periodic with the same period as the drive and describes the periodic micro-motion of the system while the Floquet exponent matrix $e^{-\mathrm{i}\mathbf{\mu}\tau}$ describes its long-term behavior.
In the supplementary material, it is shown that the reflection coefficient can be decomposed in Fourier series as:
$\tilde{r}(\tau,\omega)=\tilde{r}_{b}(\omega)+\sum_{n}\tilde{r}_{n}(\omega)e^{\mathrm{i}2n\omega_{d}\tau}$ (15)
where $\tilde{r}_{b}(\omega)$ corresponds to a non-resonant background contribution to the reflectivity and the $n^{th}$ order Fourier harmonic $\tilde{r}_{n}(\omega)$ can be expressed as:
$\frac{\tilde{r}_{n}(\omega)}{-\mathrm{i}\gamma}=\penalty 10000\ \frac{\tilde{c}_{n}(\omega-\tilde{\omega}_{d})+\tilde{d}_{n}}{(\omega-\tilde{\mu}_{+})(\omega-\tilde{\mu}_{-})}$ (16)
Here, $\gamma$ corresponds to the radiative power decay rate, the Floquet eigenvalues $\tilde{\mu}_{\pm}$ are defined through $\tilde{\mu}_{\pm}=\tilde{\omega}_{d}\pm\Delta\tilde{\mu}$ with $\tilde{\omega}_{d}\overset{\text{def}}{=}\omega_{d}-i\Gamma$ and can be obtained through diagonalization of the Floquet matrix $\mathbf{\mu}$, while the coefficients $(\tilde{c}_{n},\tilde{d}_{n})$ are related to the Fourier coefficients of the $\mathbf{G}(\tau)$ matrix. Their complete expression can be found in the supplementary material.
Spectral lineshapes (Figure 3a,d)
The experimental lineshapes of $\tilde{r}_{0}(\omega)$ and $\tilde{r}_{\pm 1}(\omega)$ are obtained from a short term Fourier transform of the time domain data (Fig. 2b) performed over a sliding window $\tau,\tau+2\pi/\omega_{d}]$ as a function of time $\tau$. They are presented in Figure 3a for $\tau\approx 24ps$. The complete temporal evolution of $\tilde{r}_{0}(\omega)$ and $\tilde{r}_{\pm 1}(\omega)$ is reported in the Supplementary material. Theoretically, they are obtained from numerical computation of formula (16).
Experimental determination of the Floquet eigenvalues (Figure 4a)
The Floquet eigenvalues presented in Fig. 4a are obtained from fitting the experimental lineshapes according to equation (16). Other coefficients are reported in the supplementary material. The fit is simultaneously performed on the real and imaginary parts of $\tilde{r}_{0}(\omega)$ and $\tilde{r}_{-1}(\omega)$ together.
As shown in the supplementary material, the fitting procedure is benchmarked on theoretical data to ensure that it accurately reproduces the eigenvalues that can obtained independently from diagonalization of the Floquet matrix.
Stability diagram (Figure 4b)